OpenStudy (anonymous):

A scientist is studying global warming. She wants to change her average temperature calculations from Celsius to Fahrenheit. The average temperature that she recorded for the Earth last year was 15 degrees Celsius. The formula for changing from Celsius to Fahrenheit is the quantity of the Celsius temperature times nine-fifths plus 32 degrees. According tot he scientist's data, what was the average temperature, in Fahrenheit, of the earth last year?

OpenStudy (tkhunny):

You could try working it out? [15ºC * (9/5)] + 32 = ?? ºF?

OpenStudy (anonymous):

So [15 * 1.8] + 32 27 + 32 = 59 F did I do that right?

OpenStudy (anonymous):

yea u did it right.i just checked it

OpenStudy (tkhunny):

Pretty easy to do in your head if you do NOT convert 9/5 to 1.8 15/5 = 3 9*3 = 27 27 + 32 = 59 Done.